Circular No. 8731

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006dp, 2006dq, 2006dr Three additional supernovae have been discovered on unfiltered CCD

images: 2006dp and 2006dr by L. A. G. Monard (cf. IAUC 8726), and 2006dq by J. Schwehr and W. Li (LOSS/KAIT; cf. IAUC 8728).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006dp July 9.12 1^{h}00^{m}38.^{s}84 − 7^{o}58′49′′.3 18.4 12′′ E, 3′′ N
2006dq July 15.31 17 59 21.01 +58 32 36.8 18.3 12′′.1 E, 9′′.4 N
2006dr July 17.15 3 17 14.19 −32 34 31.8 16.1 12′′ E, 2′′ N

Additional red magnitudes from Monard for 2006dp in MCG –01-3-56: 2005 Dec. 30.775 UT, [19.0; 2006 June 6.161, [18.2; July 11.164, 18.5 ± 0.5; 12.156, 18.3 ± 0.3. Nothing is visible at the location of 2006dp on a red image of the Digitized Sky Survey (limiting mag 20.5). SN 2006dp is a type-II supernova (spectroscopic details on CBET 575). Additional KAIT magnitudes for 2006dq in UGC 11089: June 17.40, [19.5; July 11.35, 18.3. SN 2006dq is a type-II supernova, roughly 3 weeks past explosion on July 18 (details on CBET 579). Additional red magnitudes for 2006dr in NGC 1288: July 4.139, [18.5; July 18.075, 16.1. Nothing is visible at the location of 2006dr on the Digitized Sky Survey (limiting red mag 20.5).

V2362 CYGNI S. Mazuk, R. J. Rudy, D. K. Lynch, and C. C. Venturini, Aerospace

Corporation; R. C. Puetter, University of California at San Diego; and R. B. Perry, Langley Research Center, NASA, report 0.47- to 2.5-μm spectroscopy of V2362 Cyg (cf. IAUC 8698, 8702, 8710) on June 14.6 UT using the VNIRIS spectrograph on the Lick 3-m telescope. The nova is still of low excitation, showing the signature Fe II lines as well as emission of C I, N I, and O I. He I emission lines remain comparatively weak. The reddening determined from the fluorescently excited O I lines is E(B − V) = 0.6; O I 777.4-nm is also present and displays a weak P-Cyg profile. The optical spectrum displays forbidden lines of [O I], [O II], and [N II]. FWHM of the lines was 1500 km/s. There is no evidence of dust in the nova ejecta. Magnitudes for V2362 Cyg: V = 12.4, J = 10.0, H = 10.1, K = 9.5.

COMET C/2006 M4 (SWAN) Further to IAUC 8729, R. H. McNaught corrects his tail p.a. to 155^{o}.

2006 July 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green